Use the given conditions to write a system of equations. Solve the system and find the numbers.
The sum of two numbers is 12.If one number is subtracted from the​ other, the result is -4. Find the numbers.
list the two numbers?
(use comma to separate answers)

Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

x+y=12
y=12-x
x-y=-4
----------------
x-(12-x)=-4
2x-12=-4
2x=-4+12
2x=8
x=8/2
x=4
---------------
x-y=-4
4-y=-4
-y=-4-4
y=+8
--------------
x-y=-4
4-(8)=-4
-4=-4
-------------
4,8

Answer by Alan3354(69443)

x + y = 12
x - y = -4
-------------- Add
2x = 8
x = 4
y = 8